In first case, two masses of 1 KG and 2KG are present on opposite sides{ends} of the frictionless pulley.


In second case, a mass of 1KG is on one side and 20 N force is applied on the other side of the pulley.


 What is the difference between the two situations?? PLS EXPLAIN where it matters.......


{TAKE g = 10 m/ sec2}


Grade:12

1 Answers

Spandan Pratyush, IIT Delhi
13 years ago

Dear James Bond

In the first case:

ur equations will be:-

T - 10 = a (where T is the tension & a is the acceleration of the blocks and 10 is the value of g), &

2*10 - T = 2*a

Solving which u get, a = 10/3 m/s^2 & T = 40/3 N

But in the second case:

when 20N force is applied, since there is no mass at the point of application of the force, the tension in the string becomes 20N.

So equation for the 10 KG weight becomes,

T - 10 = a, which gives a = 10 m/s^2.

?Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts

Spandan Pratyush